OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!! solve m^-5=1/32

OpenStudy (anonymous):

you change the negative power with postive

OpenStudy (anonymous):

how we change the negative 1/("the equation");

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

1/m^5 = m^-5

OpenStudy (anonymous):

\[m^-5 = 1/m^5\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/m^5 = 1/32 if you made \[\sqrt[5]{m}\]

OpenStudy (anonymous):

or you can see the 5th root for the two sides you will get th result

OpenStudy (anonymous):

i dont know how to figure out square root i havn't got that far yet

OpenStudy (anonymous):

you accept me that you have right now :- 1/m^5 = 1/32

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we have find what m is

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

the num with power 5 is 32

OpenStudy (anonymous):

the 5th root of 32 is 2; then m =2

OpenStudy (anonymous):

ok thank you